draw graph of y=sin||x-1|-1.JUST EXPLAIN ME LAST STEP THAT IS WHOLE MODULUS OF MOD (X-1)-1

Grade:12

1 Answers

Askiitians_Expert Yagyadutt
12 years ago

Hello Shivam ..! in case of modulus function ...  we try to get the value of x for which the function changes sign ..!

y = sin ( | |x-1| - 1 | )

For x>1

y = sin ( | x -2 | )

For x > 2

y = sin (x-2) and

for x <2

y = -sin(x-2)

So finally ...

for 1<x<2

y = -sin(x-2) ..................

for x>2

y = sin(x-2)

 

Now you can draw the graph here for

x > 1

 

For x < 1????

y = sin ( |-x +1 - 1| )

y = sin ( |-x| )

For x < 0 y = -sin(x)

For 0<1 y = sin (x)

 

Hence function is something like this ...

x < 0 ------------ y = -sin x

0 < x < 1 ----------- y = sin x

1 <x < 2 ------------y = -sin(x-2)

x > 2......................y = sin (x-2)

 

Graph is simple i guess ....!

 

Regards

 

Yagya

Askiitians_expert